A rock is thrown horizontally from the top of a cliff 88 m high, with a horizontal speed of 25 m/s. How far from the base of the cliff does the rock land?


1
Expert's answer
2021-05-24T18:03:24-0400

"h=\frac{gt^2}{2}\implies t=\sqrt{\frac{2h}{g}},"

"l=vt=v \sqrt{\frac{2h}{g}}=106~m."